OpenStudy (anonymous):

Mr. Doodle’s grade distribution over the past 3 years for a course in college algebra is shown in the chart below. Grade Number A 45 B 180 C 210 D 95 F 65 I 5 If Jane plans to take a college algebra course with Mr. Doodle, determine the empirical probability that she receives an A.

OpenStudy (anonymous):

there were 45 people with an A out of a total of 600 students. which makes the probability that Jane receives an A 45/600

OpenStudy (anonymous):

prob(a)/sum of all nums= 45/600=3/40